OpenStudy (anonymous):

A water truck's tank has internal dimensions of 11 meters in length and 2 meters in diameter. If the truck's load of pure water is 70 percent of full capacity, how heavy is the load in pounds? Water's density is 62.4 lb/ft^3.

OpenStudy (anonymous):

please show all calc

OpenStudy (anonymous):

Also In kilograms, what is the weight of the load if the truck carries liquid nitrogen instead of water? The liquid nitrogen's density is 0.808 g/ml.

OpenStudy (mertsj):

\[V _{water}=.7(\pi)(1^2)(11)=.77\pi m^3\]

OpenStudy (mertsj):

\[.77\pi m^3\times\frac{3.28^{3}ft ^{3}}{1m ^{3}}\times\frac{62.4 lbs}{1ft^3}=5326.6lbs\]

OpenStudy (mertsj):

\[.77\pi m ^{3}\times\frac{100^{3}ml}{1m^3}\times\frac{.808 g}{1 ml}\times\frac{1 kg}{1000 g}=1954.6 kg\]

OpenStudy (anonymous):

hey mertsj it rejected both answers

OpenStudy (mertsj):

Did you put them in by rounding?

OpenStudy (mertsj):

you might try 5327 lbs and 1955 kg

OpenStudy (mertsj):

Did you check the arithmetic?

OpenStudy (anonymous):

well im actually in class right now looking at my homework on my laptop. this is a question i have no idea of so i decided to post it. i cant check as im in class

OpenStudy (mertsj):

ok

OpenStudy (mertsj):

Ask your teacher if the calculation is set up correctly.

OpenStudy (anonymous):

ok